OpenStudy (starryknight):

Find the sum of the first 28 terms of the following sequence: -7, -4, -1 . . . 938 1876 134 588

OpenStudy (anonymous):

134

OpenStudy (anonymous):

This problem is "sequence of numbers with common difference". Suppose that d is common difference, n is a number of term. \[a_{1} = -7 = -7 +3(n-1)\] d = 3. Because the sequence increase by 3. So when n=28, \[a_{28} = -7 + 3\times (28 - 1) = -7 + 3 * 27 = 74\] Therefore, the sum of the first 28 terms is \[S_{28} = a_{1} + a_{2} +.... + a_{28} = \frac{ n (a_{1} + a_{28}) }{ 2 }\] \[=\frac{ 28 \times (-7 + 74) }{ 2 } = 14 \times 67 = 938\] The answer is "938".